1) The flow of heat (Q) through a material is given by Q = (K.A.O.)/ d
a) Transpose the formula to make A the subject.
b) find the value of A if Q = 300, O = 40 , d = 0.5 and K = 24

Answer by stanbon(75887)

1) The flow of heat (Q) through a material is given by
Q = (K*A*O)/ d
---
a) Transpose the formula to make A the subject.
KAO = dQ
A = dQ/KO
--------------------
b) find the value of A if Q = 300, O = 40 , d = 0.5 and K = 24
A = 0.5*300/24*40
A = 150/960 = 5/32
========================
Cheers,
Stan H.
================